OpenStudy (anonymous):

I HAVE TWO MORE QUESTIONS i NEED HELP WITH TO MAKE SURE THEY ARE CORRECT, PLEASE HELP:) Use the “difference of squares” rule to factor the following expression: 49x^2-4y^2? Please explain like problemsolver did, she was awesome!

OpenStudy (lalaly):

\[(a^2x^2-b^2y^2)= (ax-by)(ax+by)\] \[49x^2-4y^2=(7x-2y)(7x+2y)\]

OpenStudy (anonymous):

and thats the answer

OpenStudy (lalaly):

yup

OpenStudy (anonymous):

you got the answer by using the formula right

OpenStudy (lalaly):

yes

OpenStudy (anonymous):

THANK YOU

OpenStudy (lalaly):

yur welcome:D